2018-06-12-MEMOMIPO » History » Version 4
Rogers, Chris, 12 June 2018 16:06
1 | 1 | Long, Kenneth | h1. 2018-06-12-MEMOMIPO |
---|---|---|---|
2 | |||
3 | h2. Joint MICE Experiment Management Office and MICE International Project Office |
||
4 | |||
5 | h3. 12th June 2018; 15:30 BST |
||
6 | |||
7 | "Phone details":http://mice.iit.edu/phonebridge.html |
||
8 | |||
9 | h4. Agenda |
||
10 | |||
11 | # *Introduction: KL* |
||
12 | # *Notes of previous meeting and status of actions: All* |
||
13 | ** *PS:* Generate two-pager to close-out common fund for next PSAG. |
||
14 | ** *DR:* Continue to push towards a persistence plan for the MICE IIT WWW site. |
||
15 | ** *DR:* Coordinate migration of mail server from IIT to jiscmail. |
||
16 | # *Project Manager's report: CW* |
||
17 | # *Software and computing status: DR* |
||
18 | # *Paper status and plans: CR* |
||
19 | # *AOB* |
||
20 | |||
21 | 2 | Rogers, Chris | h4. Notes |
22 | |||
23 | Present: |
||
24 | Chris Rogers |
||
25 | Peter Garbinicius |
||
26 | Colin Whyte |
||
27 | Durga Rajaram |
||
28 | |||
29 | * Introduction - KL is in another meeting, CR to chair |
||
30 | 4 | Rogers, Chris | |
31 | 2 | Rogers, Chris | # *Notes of previous meeting and status of actions: All* |
32 | ** *PS:* Generate two-pager to close-out common fund for next PSAG. |
||
33 | *** No report - stands |
||
34 | ** *DR:* Continue to push towards a persistence plan for the MICE IIT WWW site. |
||
35 | *** Purchase order has been placed for webserver at IIT. Static mirror at RAL to be coordinated with Chris Brew. |
||
36 | ** *DR:* Coordinate migration of mail server from IIT to jiscmail. |
||
37 | *** KL will migrate EB and CB mailing lists to jiscmail |
||
38 | *** CR will migrate analysis mailing list to jiscmail |
||
39 | *** DR will migrate software mailing list to jiscmail |
||
40 | *** Other mailing lists are inactive and will be archived |
||
41 | # *Project Manager's report: CW* |
||
42 | ** See slides |
||
43 | ** Since VC |
||
44 | *** Tony will go into the synchrotron at next opportunity to check a few survey things |
||
45 | *** ISIS have offered resource to help clear MICE hall |
||
46 | *** Shipping container for US equipment now expected to arrive w/c 25th June; Alan Br likely to be available in that week |
||
47 | # *Software and computing status: DR* |
||
48 | ** V.Pec has generated a set of new TOF calibrations. TOF bad channel patch has also been brough in. DR running some tests, in time for next release. |
||
49 | ** 2017/03 geometry has wrong alignment corrections; fix is in hand. |
||
50 | ** Global track matching for backtracking from TKD is now in test. Will not make next release. |
||
51 | 3 | Rogers, Chris | ** Discussion regarding "cuts" routines; they should be run by default; DR asked CR to check the default values are reasonable |
52 | ** Handover of computing project |
||
53 | *** Archiving mostly done |
||
54 | *** Needs some one to manage any data issues and new users (for GRID) |
||
55 | *** Durga needs to hand over geeometry and "all those things". *Action: DR* to make a list. |
||
56 | 2 | Rogers, Chris | # *Paper status and plans: CR* |
57 | ** No update since VC |
||
58 | # *AOB* |
||
59 | ** No AOB |
||
60 | |||
61 | 1 | Long, Kenneth | h4. Summary of actions: |
62 | |||
63 | h3. Back to MEMO: |
||
64 | |||
65 | [[memo:|Back]] |
||
66 | |||
67 | h3. Back to MIPO: |
||
68 | |||
69 | [[mipo:|Back]] |
||
70 | |||
71 | h2. Documents |